# HybridoNet-Adapt: A Domain-Adapted Framework for Accurate Lithium-Ion Battery RUL Prediction

### v1

@@ 4.1 Experimental Setup
Our proposed RUL models are implemented using the PyTorch framework and optimized with the AdamW algorithm [27] to minimize the respective loss functions. Experiments are performed on an NVIDIA A100 GPU with 80GB of memory. Each experiment runs for 10 epochs with a batch size of 128 and a fixed learning rate of 0.0005. To obtain a representative prediction, each experiment is repeated 10 times, and the average of their predictions is taken. The training dataset is split into 90% for training and 10% for validation, with the best selected based on the lowest RMSE (see 4.3) on the validation set. The weighting factor $\lambda$ in 3.3 is dynamically adjusted during training using the following function: $\lambda=\frac{2}{1+e^{-10\cdot\frac{\text{epoch}}{\text{epochs}}}}-1.$ Here, epoch is the current training interation.
@@ 4.2.1 First dataset: Varied fast-charging conditions, with consistent discharging conditions
The first dataset, referred to as the TRI dataset [38], encompasses a detailed study of 124 LFP/graphite lithium-ion batteries. Each LIB in the dataset has a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V. The cycle life span of these batteries ranges from 150 to 2,300 cycles, showcasing a wide spectrum of longevity. The dataset is divided into three distinct parts: a training set with 41 LIBs, a primary test set with 43 LIBs, and a secondary test set comprising 40 LIBs. In terms of operational conditions, all LIBs were subjected to uniform discharge protocols. Specifically, they were discharged at a constant current rate of 4 C until the voltage dropped to 2 V, followed by a constant voltage discharge at 2 V until the current diminished to C/50. The LIBs were charged at rates between 3.6 C and 6 C, under a controlled temperature of 30°C within an environmental chamber. The dataset contains approximately 96,700 cycles, making it one of the largest datasets to consider various fast-charging protocols.
@@ 4.2.2 Second dataset: Varied discharge conditions, with consistent fast-charging conditions
The second dataset, referred as the LHP dataset [30], was developed through a battery degradation experiment involving 77 cells (LFP/graphite A123 APR18650M1A) with a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V. Each of the 77 cells was subjected to a unique multi-stage discharge protocol, while maintaining an identical fast-charging protocol for all cells. The experiment was conducted in two thermostatic chambers at a controlled temperature of 30°C. The dataset encompasses a total of 146,122 discharge cycles, making it one of the largest datasets to consider various discharge protocols. The cells exhibit a cycle life ranging from 1,100 to 2,700 cycles, with an average of 1,898 cycles and a standard deviation of 387 cycles. The discharge capacity as a function of cycle number reveals a wide distribution of cycle lives. For analytical purposes, the cells were divided into two groups: 55 cells were used for training, and 22 cells were reserved for testing. Both the observed cycle life and initial discharge capacity distributions approximate normal distributions with significant standard deviations.
@@ 4.3 Evaluation Metrics
To evaluate RUL prediction, we use Root Mean Square Error ($RMSE$) [45], R-squared ($R^{2}$) [3, 46, 43], and Mean Absolute Percentage Error ($MAPE$) [6]. These are calculated as follows:
@@
4.4 Data Analysis and Preprocessing
Figures 3(a) and 3(b) analyze battery cycle life. Figure 3(a) tracks an individual cell’s charge and discharge capacities, marking EOL when the maximum capacity degrades to 80% of nominal capacity. Figure 3(b) compares cycle life across cells, revealing significant variation in discharge capacity. This variability challenges prediction models for RUL, emphasizing the need for accurate predictions for BHM systems.
(a) Maximum charge and discharge capacities over charge-discharge cycles for a single battery cell
(b) Maximum discharge capacities over charge-discharge cycles for many battery cells in the first dataset
@@ 4.4 Data Analysis and Preprocessing
Before feature extraction, the raw signals exhibit sudden peaks and fluctuations, resembling noise. Smoothing the time-series data can help reduce noise and enhance the key characteristics of the signal. To achieve this, a median filtering method is applied to eliminate abrupt peaks in the signals before feature extraction. As a result, the application of median filtering improves overall model performance. The filtered data leads to better RMSE, $R^{2}$, and MAPE (%) values compared to the unfiltered data, as illustrated in Figure 4.
@@ 4.5 Feature Extractor
The feature extractor in both HybridoNet and HybridoNet-Adapt is progressively developed, starting with an LSTM architecture and sequentially integrating Multihead Attention (MA) and a NODE block. To evaluate the effectiveness of each component, we assess the performance of HybridoNet at different stages. With each addition, the model’s predictive capability improves. Ultimately, HybridoNet achieves an RMSE of 166.33, an $R^{2}$ score of 0.86, and a MAPE of 7.44%, demonstrating its superior performance.
Model RMSE $\downarrow$ $\uparrow$ MAPE (%) $\downarrow$
LSTM 176,66 0,83 8,35
LSTM + MA 167,01 0,85 7,53
LSTM + MA + NODE 166.33 0.86 7.44
Table 2: Performance comparison of LSTM-based models on the testing data of the second dataset.
4.6 HybridoNet-Adapt: An Extension of HybridoNet for Domain Adaption
HybridoNet-Adapt is evaluated with various feature loss functions, including CORAL Loss, Domain Loss [10], MMD, as well as combinations such as MMD with Domain Loss and MMD with Domain Loss and CORAL Loss, as shown in Figure 6. The results indicate that using only MMD as the feature loss function yields the best performance, achieving an RMSE of 160.05.
Figure 5: Comparison of different feature loss methods.
@@ 4.6 HybridoNet-Adapt: An Extension of HybridoNet for Domain Adaption
Figure 6: Comparison of RMSE across different number of LSTM layers, hidden dimensions, and dropout configurations.
To identify the optimal time step in the sequence dimension for both MA and NODE outputs, a comprehensive evaluation was conducted. Various NODE output time steps ranging from 2 to 6 were tested, along with different MA output time step selections, including the last, the second-to-last, and the mean time step. As shown in Figure 7, the best performance was achieved when using the second-to-last time step of the MA output and a NODE output time step of 2 ($t+2$).
Figure 7: Comparison of RMSE for different NODE discrete time steps ($t$) and Multihead Attention output time step selections.
The proposed HybridoNet-Adapt model is systematically evaluated under various scenarios by experimenting with four different target datasets, each derived from the training data of the second dataset. The source data is the training data from the first dataset. Below are four groups of battery cells selected from the training data of the second dataset. These groups are carefully formed to ensure each set represents a diverse range of battery performance. For instance, Group 1 includes both high-cycle cells (e.g., 2-2 with 2,651 cycles) and low-cycle cells (e.g., 1-6 with 1,143 cycles), ensuring a comprehensive representation of aging behaviors.
@@
Table 3: Comparison of HybridoNet, DANN, and HybridoNet-Adapt across four target data groups from the second dataset, based on testing scores from the testing data of the second dataset.
@@
Table 4: Evaluation metrics for RUL prediction performance on the test data from the second dataset, using existing Elastic Net, $A_{1}$, and $A_{2}$ results of paper [30] (see Table S4).
@@
Table 5 presents the performance comparison of different models on the secondary test data from the first dataset. HybridoNet-Adapt achieves the best results, with the lowest RMSE (146.52), highest $R^{2}$ score (0.72), and lowest MAPE (11.85%), demonstrating its superior predictive accuracy. HybridoNet outperforms XGBoost in all metrics, highlighting the effectiveness of deep learning-based approaches. The improvements in HybridoNet-Adapt further validate the benefits of domain adaptation in enhancing RUL prediction performance.
@@
Table 5: Evaluation metrics for RUL prediction performance on the secondary test data from the first dataset.
Figure 8 illustrates the RUL predictions of XGBoost, HybridoNet, HybridoNet-Adapt, and DANN, compared to the true (observed) RUL for Cell 4-5 and Cell 3-1 in the testing set of the second dataset. Among all methods, HybridoNet-Adapt demonstrates the closest alignment with the observed RUL, highlighting its superior predictive accuracy. This improvement is attributed to HybridoNet-Adapt’s ability to align feature representations from the source domain to the target domain, as shown in Figure 9. By effectively increasing the amount of target-relevant data through our DA adaptation technique, HybridoNet-Adapt enhances robustness, making it more adaptable to diverse real-world battery degradation scenarios.
@@
Figure 9: PCA-based comparison of embedding features between HybridoNet-Adapt and DANN.
@@ 4.7 Comparison with state-of-the-art methods
Figure 10: Comparison of our proposed models with existing state-of-the-art methods on the testing data of the second dataset.
@@ 5 Conclusion
In this study, we introduced HybridoNet, a novel deep learning framework that seamlessly integrates signal processing with LSTM, Multihead Attention, and NODE blocks to enhance RUL prediction accuracy. Additionally, we proposed HybridoNet-Adapt, an advanced variant incorporating domain adaptation inspired by DANN, significantly improving generalization across diverse battery degradation patterns. Our experimental results demonstrate the superiority of our approach over traditional machine learning models such as XGBoost and Elastic Net, as well as state-of-the-art deep learning methods like Dual-input DNN, highlighting its effectiveness in predictive maintenance. For future work, we plan to further enhance model generalization through self-supervised learning, refine real-time deployment strategies, and explore multi-modal data integration to improve scalability and robustness across various industrial applications.
Acknowledgment

### v2

@@ 4.1 Experimental Setup
Our proposed RUL model is implemented using the PyTorch framework and optimized using the AdamW algorithm [33] to minimize the respective loss functions. All experiments are conducted on an NVIDIA A100 GPU with 80GB of memory. Each experiment is trained for 10 epochs with a batch size of 128 and a fixed learning rate of 0.0005. To reduce variability in the training process, each experiment is repeated 10 times, and the final prediction is computed as the average of these runs. The training data is divided into 90% for training and 10% for validation, with the model selected based on the lowest RMSE on the validation set (see Section 4.3). The weighting factor $\lambda$ in Equation 3.2 is dynamically adjusted during training using the following schedule [13]:
$\lambda=\frac{2}{1+e^{-10\cdot\frac{\text{epoch}}{\text{epochs}}}}-1,$
where epoch denotes the current training epoch, and epochs is the total number of training epochs.
@@ 4.2.1 First dataset: Varied fast-charging conditions, with consistent discharging conditions
The first dataset, referred to as the TRI dataset [44], encompasses a detailed study of 124 LFP/graphite lithium-ion batteries. Each LIB in the dataset has a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V. The cycle life span of these batteries ranges from 150 to 2,300 cycles, showcasing a wide spectrum of longevity. In terms of operational conditions, all LIBs were subjected to uniform discharge protocols. Specifically, they were discharged at a constant current rate of 4 C until the voltage dropped to 2 V, followed by a constant voltage discharge at 2 V until the current diminished to C/50. The LIBs were charged at rates between 3.6 C and 6 C, under a controlled temperature of 30°C within an environmental chamber. The dataset contains approximately 96,700 cycles, making it one of the largest datasets to consider various fast-charging protocols. The dataset is divided into three distinct parts: a training set with 41 LIBs, a primary test set with 43 LIBs, and a secondary test set comprising 40 LIBs.
@@ 4.2.2 Second dataset: Varied discharge conditions, with consistent fast-charging conditions
The second dataset, referred as the LHP dataset [36], was developed through a battery degradation experiment involving 77 cells (LFP/graphite A123 APR18650M1A) with a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V. Each of the 77 cells was subjected to a unique multi-stage discharge protocol, while maintaining an identical fast-charging protocol for all cells. The experiment was conducted in two thermostatic chambers at a controlled temperature of 30°C. The dataset encompasses a total of 146,122 discharge cycles, making it one of the largest datasets to consider various discharge protocols. The cells exhibit a cycle life ranging from 1,100 to 2,700 cycles, with an average of 1,898 cycles and a standard deviation of 387 cycles. The discharge capacity as a function of cycle number reveals a wide distribution of cycle lives. The dataset is divided into two distinct parts: a training set with 55 LIBs, and a test set with 22 LIBs.
@@ 4.3 Evaluation Metrics
To evaluate RUL prediction, we use Root Mean Square Error ($RMSE$) [50], R-squared ($R^{2}$) [5, 49], and Mean Absolute Percentage Error ($MAPE$) [9]. These are calculated as follows:
@@
4.4 Signal Analysis and Preprocessing
Figures 3(a) and 3(b) analyze battery cycle life. Figure 3(a) tracks an individual cell’s charge and discharge capacities, marking EOL when the maximum capacity degrades to 80% of nominal capacity. Figure 3(b) compares cycle life across cells, revealing significant variation in discharge capacity. This variability challenges prediction models for RUL, emphasizing the need for accurate and adaptable RUL predictions for BHM systems.
(a) Maximum charge and discharge capacities over charge-discharge cycles for a single battery cell.
(b) Maximum discharge capacities over charge-discharge cycles for many battery cells.
@@ 4.4 Signal Analysis and Preprocessing
Before feature extraction step in the signal preprocessing phase (as mentioned in 3), the raw signals exhibit sudden peaks and fluctuations, resembling noise. Smoothing the time-series data can help reduce noise and enhance the key characteristics of the signal. To achieve this, a median filtering method is applied to eliminate abrupt peaks in the signals before feature extraction. As a result, the application of median filtering improves overall model performance. The filtered data leads to better RMSE, $R^{2}$, and MAPE (%) values compared to the unfiltered data, as illustrated in Figure 4.
@@ 4.5 Feature Extractor
The feature extractor is progressively developed, starting with an LSTM architecture and sequentially integrating Multihead Attention (MA) and a NODE block. To evaluate the effectiveness of each component, we assess the performance of HybridoNet-Adapt at different stages. With each addition as shown in Figure 5, the model’s predictive capability improves. Ultimately, HybridoNet-Adapt achieves an RMSE of 166.33, an $R^{2}$ score of 0.86, and a MAPE of 7.44%, demonstrating its superior performance.
Figure 5: Performance comparison of LSTM-based blocks for feature extractor. Experiment on the testing data of the second dataset.
4.6 HybridoNet-Adapt with Domain Adaption
HybridoNet-Adapt is evaluated with various feature loss functions, including CORAL Loss, Domain Loss [14], MMD, as well as combinations such as MMD with Domain Loss, and MMD with Domain Loss and CORAL Loss, as shown in Figure 6. The results indicate that using only MMD as the feature loss function yields the best performance, achieving an RMSE of 160.05.
Figure 6: Comparison of different feature loss methods. Experiment on the testing data of the second dataset.
@@ 4.6 HybridoNet-Adapt with Domain Adaption
Figure 7: Comparison of RMSE across different number of LSTM layers, hidden dimensions, and dropout configurations. Experiment on the testing data of the second dataset.
To identify the optimal time step in the sequence dimension for both Multihead Attention and NODE outputs, a comprehensive evaluation was conducted. Various NODE output time steps ranging from 2 to 6 were tested, along with different Multihead Attention output time step selections, including the last, the second-to-last, and the mean time step. As shown in Figure 8, the best performance was achieved when using the second-to-last time step of the Multihead Attention output and a NODE output time step of 2.
Figure 8: Comparison of RMSE for different NODE discrete time steps ($t$) and Multihead Attention output time step selections. Experiment on the testing data of the second dataset.
The proposed HybridoNet-Adapt model is systematically evaluated under various scenarios by experimenting with four different target sets, each derived from the training data of the second dataset. The source data is the training data from the first dataset. Below are four target groups of battery cells selected from the training data of the second dataset. These groups are carefully formed to ensure each set represents a diverse range of battery performance. For instance, Group 1 includes both high-cycle cells (e.g., 2-2 with 2,651 cycles) and low-cycle cells (e.g., 1-6 with 1,143 cycles), ensuring a comprehensive representation of aging behaviors.
@@
Table 2: Comparison of HybridoNet, DANN, and HybridoNet-Adapt across four target data groups from the testing data of the second dataset.
@@
Table 3: Evaluation metrics for RUL prediction performance using existing Elastic Net, $A_{1}$, and $A_{2}$ results of paper [36] (see Table S4). Experiment on the testing data of the second dataset.
@@
Table 4 presents a performance comparison of different models on the secondary test data from the first dataset. XGBoost and HybridoNet are trained using the training data from the first dataset. HybridoNet-Adapt is trained with the second dataset’s training data as the source input, and the first dataset’s training data as the target input. HybridoNet-Adapt achieves the best results, with the lowest RMSE (146.52), the highest $R^{2}$ score (0.72), and the lowest MAPE (11.85%), demonstrating its superior predictive accuracy with domain adaptation. Additionally, HybridoNet outperforms XGBoost in all metrics, highlighting the effectiveness of deep learning-based approaches. The improvements seen in HybridoNet-Adapt further validate the benefits of domain adaptation in enhancing RUL prediction performance.
Table 4: Evaluation metrics for RUL prediction performance. Experiment on the secondary testing data of the first dataset.
@@
Figure 9 illustrates the RUL predictions of XGBoost, HybridoNet, HybridoNet-Adapt, and DANN, compared to the true (observed) RUL for Cell 4-5 and Cell 3-1 in the testing set of the second dataset. Among all methods, HybridoNet-Adapt demonstrates the closest alignment with the observed RUL, highlighting its superior predictive accuracy. This improvement is attributed to HybridoNet-Adapt’s ability to align feature representations from the source domain to the target domain, as shown in Figure 10. By effectively increasing the amount of target-relevant data through our domain adaptation technique, HybridoNet-Adapt enhances robustness, making it more adaptable to diverse real-world battery degradation scenarios.
@@
Figure 10: PCA-based comparison of embedding features between HybridoNet-Adapt and DANN from Cell 3-1 the testing data of the second dataset.
@@ 4.7 Comparison with state-of-the-art methods
Figure 11: Comparison of our proposed models with existing state-of-the-art methods. Experiment on the testing data of the second dataset..
@@ 5 Conclusion
In this study, we proposed a Remaining Useful Life (RUL) prediction framework comprising a signal processing module and a novel prediction model trained using a domain adaptation technique. In the signal preprocessing phase, raw signals undergo a noise-reduction filter, followed by feature extraction and normalization to prepare the data for prediction model input. The proposed prediction model, HybridoNet-Adapt, consists of a feature extraction block and two predictors, which are constructed using LSTM layers, Multihead Attention mechanisms, Neural Ordinary Differential Equation (NODE) blocks, and linear layers. The outputs of the two predictors are balanced using two trainable trade-off parameters. To train the model, we introduce a domain adaptation strategy that combines Mean Squared Error (MSE) with Maximum Mean Discrepancy (MMD), enabling the model to learn domain-invariant features from the source domain and effectively generalize to the target domain. Experimental results demonstrate that HybridoNet-Adapt outperforms conventional machine learning models such as XGBoost and Elastic Net, as well as state-of-the-art deep learning models like Dual-input DNN, affirming its effectiveness in the RUL prediction task. For future work, we plan to enhance model generalization through self-supervised learning, optimize strategies for real-time deployment, and explore multi-modal data integration to improve scalability and robustness across diverse battery health management (BHM) applications.
Preprint Availability
A preprint of this manuscript is available at: https://arxiv.org/pdf/2503.21392